News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 6 September 1999 9.00am EST

The LETG retract via hardware command was completed successfully at 8AM 9/6. The LETG was then inserted and retracted via software command, and was also successful. Prior to the software commanding, a relay was switched to isolate the LETG B-side windings of the MCE. As expected, zero volts were seen from the B-side, confirming that the source of the approx 6 volts seen on the B-side during insert/retract is associated with back-EMF. Analysis will continue to confirm that the observed value and behavior is fully understood.

During the night the observation of Capella was conducted and the team confirmed the expected count rate in the zero order and confirmed that Fe XVII lines were visible at approx 15-17 A.





Return to Status Reports